OpenStudy (anonymous):

when monochromatic light of wavelength 5500x10^-10 shines on a metal , photoelectrons with maximum kinetic energy 1.02eV are emitted . if the wavelength of the incident light changes to 4800x10^-10 , the max kinetic energy of photoelectrons becomes 1.35eV . calculate the value of Planck constant based on information given , can somebody try solve this question and explain for me , i don't get it all . please.

OpenStudy (anonymous):

this sum involves use of planks theorey and photoelectric effect by planck's theorey E=h*(frequency) where E is the energy h is planck's constant frequency=c/wavelength where c is speed of light there is a minmum frequency(threshold frequency needed for ejection of an electron) E0 gives the energy associared with it E=E0+K.E k.e is kinetic energy so E0=E-K.E===> 1 E0 is going to remain the same whatever the case is find E1 and then E2 finding frequencies in E=h*frequency ,leave h alone use equation 1 and equate both equations and find h be back when u have doubt in this....

OpenStudy (anonymous):

hc/lemda1=work function+(maximum kinetic energy)1..(1) hc/lemda2=work function+(maximum kinetic energy)2...(ii) lemda1=5500*10^-10 m,(maximum kinetic energy)1=1.02ev lembda2=4800x10^-10m,(maximum kinetic energy)2=1.35ev c=3*10^8m/sec using this two equation ,find the value 0f h.